# Adaptive Off-Policy Inference for M-Estimators Under Model Misspecification

### v1

@@ 4 Practical Strategies for Covariance Estimation
Our strategy for building an estimator of the variance is to decompose $V_{t,\theta^{\star}}$ into terms that are independent of the history and then find empirical estimators of these quantities from external data, similar in spirit to Kato (2020).
@@
$\mathrm{Var}(s_{t,\theta^{\star}}\mid\mathcal{H}_{t-1})=\mathrm{Var}\Big[\mathbb{E}_{A_{t}\sim\pi_{e}}\big[\dot{m}_{\theta^{\star}}(X_{t},A_{t},Y_{t})\mid X_{t}\big]\Big]+\mathbb{E}\Big[\mathrm{Var}_{A_{t}\sim\tilde{\pi}_{e,t}}\big(\dot{m}_{\theta^{\star}}(X_{t},A_{t},Y_{t}(A_{t}))\mid X_{t}\big)\;\Big|\;\mathcal{H}_{t-1}\Big],$
where $\tilde{\pi}_{e,t}(a\mid X_{t}):=\frac{\pi_{e}(a\mid X_{t})^{2}}{\mathbb{P}(A_{t}=a\mid X_{t},\mathcal{H}_{t-1})}$.
@@ 4.1 Using external data
There exists an external data set $\tilde{X}:=\{\tilde{X}_{i}\}_{i=1}^{n}$, independent of $\mathcal{H}_{t}$ for all $t\in[T]$, where $\tilde{X}_{i}\overset{\mathrm{iid}}{\sim}p(x)$ and $n=[rT]$ for some fixed $r\in(0,\infty)$.
@@
We can construct functions $g_{t}:\mathbb{R}\times\mathcal{A}\to\mathbb{R}^{d}$ and $h_{t}:\mathbb{R}\times\mathcal{A}\to\mathbb{R}^{d\times d}$ adapted to $\sigma(\mathcal{H}_{t-1})$ such that $g_{t}(X_{t},a)-\mathbb{E}\left[\dot{m}_{\theta^{\star}}(X_{t},A_{t},Y_{t})|X_{t},A_{t}=a\right]\overset{p}{\to}0$ for all $a\in\mathcal{A}$ and $h_{t}(X_{t},a)-\mathbb{E}\left[\dot{m}_{\theta^{\star}}(X_{t},A_{t},Y_{t})\dot{m}_{\theta^{\star}}(X_{t},A_{t},Y_{t})^{T}|X_{t},A_{t}=a\right]\overset{p}{\to}0$ for all $a\in\mathcal{A}$.
We postpone discussion of how to construct the functions $g_{t}$ and $h_{t}$ to Proposition 2. For now, we take the existence of these quantities as given and use them to construct an estimate for $\mathrm{Var}\left(s_{t,\theta^{\star}}|\mathcal{H}_{t-1}\right)$.
@@
$\hat{V}_{t}$ $=\frac{1}{n_{r}-1}\Biggl(\sum_{X_{i}\in\mathcal{H}_{t}^{(2)}}\hat{\nu}_{t}(X_{i})^{2}-\frac{1}{n_{r}}\sum_{X_{i}\in\mathcal{H}_{t}^{(2)}}\hat{\nu}_{t}(X_{i})\Biggr)\Biggl(\sum_{X_{i}\in\mathcal{H}_{t}^{(2)}}\hat{\nu}_{t}(X_{i})^{2}-\frac{1}{n_{r}}\sum_{X_{i}\in\mathcal{H}_{t}^{(2)}}\hat{\nu}_{t}(X_{i})\Biggr)^{T}$
$\quad+\frac{1}{n_{r}}\sum_{X_{i}\in\mathcal{H}_{t}^{(2)}}\sum_{a\in\mathcal{A}}\Biggl[\frac{\pi_{e}(A_{t}=a\mid X_{i})^{2}\,h_{t}(X_{i},a)}{\mathbb{P}(A_{t}=a\mid X_{i},\mathcal{H}_{t-1}^{(2)})}-\pi_{e}(A_{t}=a\mid X_{i})^{2}\,g_{t}(X_{i},a)\Biggr].$
@@
Assume you have access to an estimator $\bar{\theta}_{T}\in\sigma(\mathcal{H}_{T})$ such that $\bar{\theta}_{T}\overset{p}{\to}\theta^{\star}$. Assume you can construct $f_{t}:\mathbb{R}\times\mathcal{A}\to\mathbb{R}$ and $e_{t}:\mathbb{R}\times\mathcal{A}\to\mathbb{R}$ adapted to $\sigma(\mathcal{H}_{t-1})$ such that $f_{t}(a,X_{t})-\mathbb{E}\left[Y_{t}|X_{t},A_{t}=a\right]\overset{p}{\to}0$ and $e_{t}(a,X_{t})-\mathbb{E}\left[Y_{t}^{2}|X_{t},A_{t}=a\right]\overset{p}{\to}0$. Furthermore, assume that $\dot{m}_{\theta}$ is linear in $y$ such that it can be decomposed as $\dot{m}_{\theta}(x,a,y)=z_{\theta}(x,a)y+\nu_{\theta}(x,a)$ for some continuous functions $z,\nu$. Then, defining
$g_{t}(a,x)=z_{\bar{\theta}_{T}}(x,a)f_{t}(x,a)+\nu_{\hat{\theta}_{T}}(x,a)$
$h_{t}(x,a)=z_{\bar{\theta}_{T}}(x,a)z_{\bar{\theta}_{T}}(x,a)^{T}e_{t}(x,a),$
@@
Although this may seem restrictive, GLMs are a notable class of models that have a score function that satisfies these conditions. In our settings, GLMs will have a score function of the form.
@@
where $z:\mathbb{R}^{d}\times|\mathcal{A}|\rightarrow\mathbb{R}^{p}$ denotes some deterministic transformations of $(X{t},A_{t})$ (e.g. one-hot encoding of $A_{t}$, interaction terms between $A_{t}$ and $X_{t}$) and $\psi(\cdot)$ is the inverse-link function that maps $\theta^{T}z(X_{t},A_{t})$ to the mean under the (possibly misspecified) working model. In this case, we have:
@@
$h_{t}(X_{t},a)=z(X_{t},a)z(X_{t},a)^{T}e_{t}(X_{t},a).$
Proposition 3 simplifies the variance estimation task, as it converts the problem of estimation of potentially very high-dimensional covariance matrices into a simpler question of estimating conditional means and variances of $Y_{t}$. Note that $f_{t}$ can be the same model that was used in the definition of the MAIPWM estimator (Equation 5) so this procedure only requires the management of an additional second model targeting the squared residuals, $(Y_{t}-f_{t})^{2}$. In the case where $m_{\theta}$ is not linear in $y$, then Jensen’s inequality implies that a straightforward plug-in estimate is not available since $\mathbb{E}\left[\dot{m}_{\theta}(X_{t},A_{t},Y_{t})\right]\neq\dot{m}_{\theta}(X_{t},A_{t},\mathbb{E}[Y_{t}])$. In these cases, more complicated methods of density estimation will be required.
Some care must also be taken when choosing $\bar{\theta}_{T}$ in Proposition 3. Although $\hat{\theta}_{T}$ is consistent, it is defined as $\frac{1}{T}\hat{V}_{t}^{-1/2}s_{t,\hat{\theta}}=o_{p}(1)$ and therefore assumes that $\hat{V}_{t}$ has already been constructed in its definition, so it is not available when we need to construct $\hat{V}_{t}$. Fortunately, $\tilde{\theta}_{T}$ can be shown to be consistent.
@@
Let $\tilde{\theta_{T}}$ be defined as in Equation 5. Under the conditions of Theorem 1, $\left\lVert\tilde{\theta}_{T}-\theta^{\star}\right\rVert_{1}=o_{p}(1)$
@@
1:Data stream $\{(X_{t},A_{t},Y_{t})\}_{t=1}^{T}$; target policy $\pi_{e}(a\mid X_{t})$; model class $\{m_{\theta}:\theta\in\Theta\}$ for the GLM with transformation function $z(X,a)$ and link function $\psi(\cdot)$.
@@
3:  Construct predictive models $f_{t},e_{t}$ using $\mathcal{H}_{t-1}$ targeting the first and second moments:
$f_{t}(X_{i},a)\approx E[Y_{i}\mid X_{i},A_{i}=a],\quad e_{t}(X_{i},a)\approx E[Y_{i}^{2}\mid X_{i},A_{i}=a].$
@@
$h_{t}(x,a)$ $\leftarrow z(x,a)z(x,a)^{T}e_{t}(x,a).$
7:Use Proposition 2 to construct $\hat{V}_{t}$ for every $t\in[T]$.
8:Use $\{\hat{V}_{t}\}_{t=1}^{T}$ and apply Theorem 1 to construct confidence intervals.
@@ 4.2 Using sequential sample splitting
Proceeding in this fashion constructs two separate histories with the same distribution that are independent of each other. Note that for this particular application, we are only trying to sample from the marginal distribution of $X_{t}$ so the assignment of $A_{t}$ is only required for one of the two histories. An alternative experimental setup might only track the pairs $(X_{t},A_{t},Y_{t})$ for $\mathcal{H}_{t}^{(1)}$ and then refrain from treating or tracking outcomes for individuals in $\mathcal{H}_{t}^{(2)}$. In contexts where there is a cost to administering the treatment, the latter design will be preferable, but in other situations it may not be possible or desirable to refrain from treatment in which case the former is required.
@@
$\zeta_{t}\sim\text{Ber}(p)$ for some fixed $r\in(0,1)$ and is independent of $\mathcal{H}_{t-1}^{(1)}$ and $\mathcal{H}_{t-1}^{(2)}$ for all $t\in[T]$.
@@
There exist functions $f_{t}:\mathbb{R}\times\mathcal{A}\to\mathbb{R}$ and $g_{t}:\mathbb{R}\times\mathcal{A}\to\mathbb{R}^{d\times d}$ adapted to $\sigma(\mathcal{H}_{t-1}^{(1)})$ such that $g_{t}(a,X_{t})-\mathbb{E}[Y_{t}|X_{t},A_{t}=a]\overset{p}{\to}0$ for all $a\in\mathcal{A}$ and $h_{t}(a,X_{t})-\mathrm{Var}\left(\dot{m}_{\theta^{\star}}(X_{t},a,Y_{t}(a))|X_{t}\right)\overset{p}{\to}0$ for all $a\in\mathcal{A}$.
11 is almost the same as 9 but we require that the models are only trained on $\mathcal{H}_{t-1}^{(1)}$ to ensure that $\mathcal{H}_{t-1}^{(2)}$ can be preserved as an independent data set for sampling features. At this point, the procedure can proceed the same as in Section 4. For completeness, we present a slightly modified version of Proposition 2 adapted to the case of sample splitting.
@@
Under Assumptions 1- 6, 10, and $1$, define $\hat{\nu}_{t}(X_{i})$ as in Equation 8 and $n_{r}:=\sum_{t=1}^{T}\zeta_{t}$. Define:
$\hat{V}_{t}$ $=\frac{1}{n_{r}-1}\Biggl(\sum_{X_{i}\in\mathcal{H}_{t}^{(2)}}\hat{\nu}_{t}(X_{i})^{2}-\frac{1}{n_{r}}\sum_{X_{i}\in\mathcal{H}_{t}^{(2)}}\hat{\nu}_{t}(X_{i})\Biggr)\Biggl(\sum_{X_{i}\in\mathcal{H}_{t}^{(2)}}\hat{\nu}_{t}(X_{i})^{2}-\frac{1}{n_{r}}\sum_{X_{i}\in\mathcal{H}_{t}^{(2)}}\hat{\nu}_{t}(X_{i})\Biggr)^{T}$
$\quad+\frac{1}{n_{r}}\sum_{X_{i}\in\mathcal{H}_{t}^{(2)}}\sum_{a\in\mathcal{A}}\Biggl[\frac{\pi_{e}(A_{t}=a\mid X_{i})^{2}\,h_{t}(X_{i},a)}{\mathbb{P}(A_{t}=a\mid X_{i},\mathcal{H}_{t-1}^{(2)})}-\pi_{e}(A_{t}=a\mid X_{i})^{2}\,g_{t}(X_{i},a)\Biggr].$

### v2

@@ 4 Practical Strategies for Covariance Estimation
Our strategy for building an estimator of the variance is to decompose $\mathrm{Var}(s_{t,\theta^{\star}}\mid\mathcal{H}_{t-1})$ into terms that are either independent of the history or known to the experimenter and then find empirical estimators of these quantities from external data, similar in spirit to Kato (2020).
@@
$\mathrm{Var}(s_{t,\theta^{\star}}\mid\mathcal{H}_{t-1})$ $=\mathrm{Var}\Big(\mathbb{E}_{A_{t}\sim\pi_{e}}\left[\dot{m}_{\theta^{\star}}(X_{t},A_{t},Y_{t})\mid X_{t}\right]\Big)+$
$\qquad\mathbb{E}\Bigg[\sum_{a=1}^{K}\frac{\pi_{e}(a\mid X_{t})^{2}}{\pi_{t}(a|X_{t})}\mathbb{E}[\dot{m}_{\theta^{\star}}(X_{t},a,Y_{t}(a))\dot{m}_{\theta^{\star}}(X_{t},a,Y_{t}(a))^{T}\mid X_{t}]\mid\mathcal{H}_{t-1}\Bigg]$-
$\qquad\mathbb{E}\left[\mathbb{E}_{A_{t}\sim\pi_{e}}\left[\dot{m}_{\theta^{\star}}(X_{t},A_{t},Y_{t})\mid X_{t}\right]\mathbb{E}_{A_{t}\sim\pi_{e}}\left[\dot{m}_{\theta^{\star}}(X_{t},A_{t},Y_{t})\mid X_{t}\right]^{T}\right].$
@@ 4.1 Using external data
There exists an external data set $\tilde{X}:=\{\tilde{X}_{i}\}_{i=1}^{n}$, independent of $\mathcal{H}_{t}$ for all $t\in T$, where $\tilde{X}_{i}\overset{\mathrm{iid}}{\sim}p(x)$ and $n=[rT]$ for some fixed $r\in(0,\infty)$.
@@
We can construct functions $g_{t}:\mathbb{R}^{p}\times\mathcal{A}\to\mathbb{R}^{d}$ and $h_{t}:\mathbb{R}^{p}\times\mathcal{A}\to\mathbb{R}^{d\times d}$ adapted to $\sigma(\mathcal{H}_{t-1})$ such that $g_{t}(X_{t},a)-\mathbb{E}\left[\dot{m}_{\theta^{\star}}(X_{t},A_{t},Y_{t})|X_{t},A_{t}=a\right]\overset{p}{\to}0$ for all $a\in\mathcal{A}$ and
$h_{t}(X_{t},a)-\mathbb{E}\left[\dot{m}_{\theta^{\star}}(X_{t},A_{t},Y_{t})\dot{m}_{\theta^{\star}}(X_{t},A_{t},Y_{t})^{T}|X_{t},A_{t}=a\right]\overset{p}{\to}0$
for all $a\in\mathcal{A}$.
We postpone discussion of how to construct the functions $g_{t}$ and $h_{t}$ to Proposition 3. In this section, we assume the existence of these quantities and use them to construct an estimate for $\mathrm{Var}\left(s_{t,\theta^{\star}}|\mathcal{H}_{t-1}\right)$.
@@
$\hat{V}_{t}$ $=\frac{1}{n-1}\Biggl(\sum_{X_{i}\in\mathcal{\tilde{X}}}\hat{\nu}_{t}(X_{i})-\bar{\nu}\Biggr)\Biggl(\sum_{X_{i}\in\mathcal{\tilde{X}}}\hat{\nu}_{t}(X_{i})-\bar{\nu}\Biggr)^{T}$
$\quad+\frac{1}{n}\sum_{X_{i}\in\mathcal{\tilde{X}}}\sum_{a=1}^{K}\frac{\pi_{e}(A_{t}=a\mid X_{i})^{2}\,h_{t}(X_{i},a)}{\mathbb{P}(A_{t}=a\mid X_{i},\mathcal{H}_{t-1})}-\frac{1}{n}\sum_{X_{i}\in\mathcal{\tilde{X}}}\hat{\nu}_{t}(X_{i})\hat{\nu}_{t}(X_{i})^{T}.$
@@
Assume you have access to an estimator $\bar{\theta}_{T}\in\sigma(\mathcal{H}_{T})$ such that $\bar{\theta}_{T}\overset{p}{\to}\theta^{\star}$. Assume that you can construct a sequence of continuous functions $f_{t}:\mathbb{R}\times\mathcal{A}\to\mathbb{R}$ and $j_{t}:\mathbb{R}\times\mathcal{A}\to\mathbb{R}$ adapted to $\sigma(\mathcal{H}_{t-1})$ such that $f_{t}(X_{t},a)-\mathbb{E}\left[Y_{t}|X_{t},A_{t}=a\right]\overset{p}{\to}0$ and $j_{t}(X_{t},a)-\text{Var}\left[Y_{t}|X_{t},A_{t}=a\right]\overset{p}{\to}0$. Furthermore, assume that $\dot{m}_{\theta}$ is linear in $y$ such that it can be decomposed as $\dot{m}_{\theta}(x,a,y)=z_{\theta}(x,a)y+\nu_{\theta}(x,a)$ for some continuous functions $z,\nu$. Then, defining
$g_{t}(x,a)=z_{\bar{\theta}_{T}}(x,a)f_{t}(x,a)+\nu_{\bar{\theta}_{T}}(x,a)$
$h_{t}(x,a)=z_{\bar{\theta}_{T}}(x,a)z_{\bar{\theta}_{T}}(x,a)^{T}j_{t}(x,a),$
@@
Although this may seem restrictive, GLMs are a notable class of models that have a score function that satisfies these conditions. In our settings, GLMs will satisfy a first order condition of the form
@@
where $z:\mathbb{R}^{d}\times|\mathcal{A}|\rightarrow\mathbb{R}^{p}$ denotes some deterministic transformations of $(X_{t},A_{t})$ (e.g., one-hot encoding of $A_{t}$, interaction terms between $A_{t}$ and $X_{t}$) and $\psi(\cdot)$ is the inverse-link function that maps $\theta^{T}z(X_{t},A_{t})$ to the mean under the (possibly misspecified) working model. In this case, we have:
@@
$h_{t}(X_{t},a)=z(X_{t},a)z(X_{t},a)^{T}j_{t}(X_{t},a).$
Proposition 3 simplifies the variance estimation task, as it converts the problem of estimation of potentially very high-dimensional covariance matrices into a simpler question of estimating conditional means and variances of $Y_{t}$. Note that $f_{t}$ can be the same model that was used in the definition of the MAIPWM estimator (Equation 6) so this procedure only requires the management of an additional second model targeting $\mathbb{E}\left[Y_{t}^{2}|X_{t},A_{t}=a\right]$. In the case where $m_{\theta}$ is not linear in $y$, then Jensen’s inequality implies that a straightforward plug-in estimate is not available since $\mathbb{E}\left[\dot{m}_{\theta}(X_{t},A_{t},Y_{t})\right]\neq\dot{m}_{\theta}(X_{t},A_{t},\mathbb{E}[Y_{t}])$. In these cases, more complicated methods of density estimation will be required.
Some care must also be taken when choosing $\bar{\theta}_{T}$ in Proposition 3. Although $\hat{\theta}_{T}$ is consistent, it is defined as an estimator that satisfies $\frac{1}{T}\sum_{t=1}^{T}\hat{V}_{t}^{-1/2}s_{t,\hat{\theta}}=o_{p}(1)$ and therefore assumes that $\hat{V}_{t}$ has already been constructed in its definition, so it is not available when we need to construct $\hat{V}_{t}$. Fortunately, $\tilde{\theta}_{T}$ can be shown to be consistent.
@@
Let $\tilde{\theta}_{T}$ be defined as in Equation 6. Under the conditions of Theorem 1, $\left\lVert\tilde{\theta}_{T}-\theta^{\star}\right\rVert_{1}=o_{p}(1)$.
@@
1:Data $\{(X_{t},A_{t},Y_{t})\}_{t=1}^{T}$; evaluation dataset $\tilde{X}:=\{\tilde{X}_{i}\}_{i=1}^{n}$; target policy $\pi_{e}(a\mid X_{t})$; model class $\{m_{\theta}:\theta\in\Theta\}$ for the GLM with transformation function $z(x,a)$ and link function $\psi(\cdot)$.
@@
3:  Construct predictive models $f_{t},j_{t}$ using $\mathcal{H}_{t-1}$ targeting the conditional mean and variances:
$f_{t}(X_{i},a)\approx E[Y_{i}\mid X_{i},A_{i}=a],\quad j_{t}(X_{i},a)\approx\text{Var}[Y_{i}\mid X_{i},A_{i}=a].$
@@
$h_{t}(x,a)$ $\leftarrow z(x,a)z(x,a)^{T}j_{t}(x,a).$
7:Use $\tilde{X}$ and Proposition 2 to construct $\hat{V}_{t}$ for every $t\in[T]$.
8:Use $\{\hat{V}_{t}\}_{t=1}^{T}$ and apply Theorem 1 to construct $\hat{\theta}_{T}$ and corresponding confidence intervals.
@@ 4.2 Using sequential sample splitting
Proceeding in this fashion constructs two separate histories with the same distribution that are independent of each other. Note that for this particular application, we are only trying to sample from the marginal distribution of $X_{t}$ so the assignment of $A_{t}$ is only required for one of the two histories. An alternative experimental setup might only track the triples $(X_{t},A_{t},Y_{t})$ for $\mathcal{H}_{t}^{(1)}$ and then refrain from treating or tracking outcomes for individuals in $\mathcal{H}_{t}^{(2)}$. In contexts where there is a cost to administering the treatment, the latter design will be preferable, but in other situations it may not be possible or desirable to refrain from treatment in which case the former is required.
@@
$\zeta_{t}\sim\text{Ber}(p)$ for some fixed $p\in(0,1)$ and is independent of $\mathcal{H}_{t-1}^{(1)}$ and $\mathcal{H}_{t-1}^{(2)}$ for all $t\in[T]$.
@@
There exist functions $g_{t}:\mathbb{R}^{p}\times\mathcal{A}\to\mathbb{R}$ and $h_{t}:\mathbb{R}^{p}\times\mathcal{A}\to\mathbb{R}^{d\times d}$ adapted to $\sigma(\mathcal{H}_{t-1}^{(1)})$ such that $g_{t}(X_{t},a)-\mathbb{E}[Y_{t}|X_{t},A_{t}=a]\overset{p}{\to}0$ for all $a\in\mathcal{A}$ and
$h_{t}(X_{t},a)-\mathbb{E}\left[\dot{m}_{\theta^{\star}}(X_{t},A_{t},Y_{t})\dot{m}_{\theta^{\star}}(X_{t},A_{t},Y_{t})^{T}|X_{t},A_{t}=a\right]\overset{p}{\to}0$
for all $a\in\mathcal{A}$.
11 is almost the same as 9 but we require that the models are only trained on $\mathcal{H}_{t-1}^{(1)}$ to ensure that $\mathcal{H}_{t-1}^{(2)}$ can be preserved as an independent data set for calculating $\hat{V}_{t}^{-1/2}$. At this point, the procedure can proceed the same as in Section 4.1. For completeness, we present a slightly modified version of Proposition 2 adapted to the case of sample splitting.
@@
Under Assumptions 1- 6, 10, and $1$, define $\hat{\nu}_{t}(X_{i})$ as in Equation 9 and $n_{p}:=\sum_{t=1}^{T}\zeta_{t}$. Define:
$\hat{V}_{t}$ $=\frac{1}{n_{p}-1}\Biggl(\sum_{X_{i}\in\mathcal{H}_{t}^{(2)}}\hat{\nu}_{t}(X_{i})-\frac{1}{n_{p}}\sum_{X_{i}\in\mathcal{H}_{t}^{(2)}}\hat{\nu}_{t}(X_{i})\Biggr)\Biggl(\sum_{X_{i}\in\mathcal{H}_{t}^{(2)}}\hat{\nu}_{t}(X_{i})-\frac{1}{n_{p}}\sum_{X_{i}\in\mathcal{H}_{t}^{(2)}}\hat{\nu}_{t}(X_{i})\Biggr)^{T}+$
$\qquad\frac{1}{n_{p}}\sum_{X_{i}\in\mathcal{H}_{t}^{(2)}}\sum_{a=1}^{K}\Biggl[\frac{\pi_{e}(A_{t}=a\mid X_{i})^{2}\,h_{t}(X_{i},a)}{\mathbb{P}(A_{t}=a\mid X_{i},\mathcal{H}_{t-1}^{(2)})}\Biggr]-\frac{1}{n_{p}}\sum_{X_{i}\in\mathcal{H}_{t}^{(2)}}\hat{\nu}_{t}(X_{i})\hat{\nu}_{t}(X_{i})^{T}.$